# GoalLadder: Incremental Goal Discovery with Vision-Language Models

### v1

@@ 4.3 How sample-efficient is GoalLadder?
Using large pretrained models is expensive, so methods that rely on them for feedback must do so efficiently. While embedding-based approaches, like VLM-RM [16], use all environment observations for reward calculation, preference-based methods, like RL-VLM-F [11], can offer better sample efficiency by training a reward function from a limited amount of feedback, in the hope that the learned reward function generalises to unseen states. By contrast, GoalLadder produces reward signals that generalise to unseen states by defining the reward in terms of distances in an embedding space (Section 3.2.3) trained with unsupervised learning. Utilising the full extent of agent’s experiences, rather than just the labelled observations, for reward definition allows GoalLadder to approximate state utility without collecting large amounts of feedback. Averaged across all Metaworld tasks, GoalLadder learns to solve them after only $\sim$4500 queries. For comparison, PEBBLE [30], a preference-based method that uses ground-truth reward preferences, achieves the same performance after attaining an average of $\sim$15000 preference labels on the same tasks. Indeed, we observe that GoalLadder boasts superior sample efficiency compared to the tested baselines. Performance curves in Figure 4 illustrate that RL-VLM-F struggles to learn the tasks using the feedback rate of GoalLadder, while VLM-RM, that embeds all collected observations, fails almost entirely.
@@ A.1 Soft-Actor Critic
The implementation of the Soft-Actor Critic (SAC) [29] used in the paper follows the implementation of SAC in PEBBLE [30], similar to Wang et al. [11] to ensure a fair comparison. As such, the used hyperparameters, including the used optimisers, are the same and can be found in the original paper [30].
@@ A.2 Feature extractor
As mentioned in the paper, we train a visual feature extractor $\psi(\cdot)$ using a variational autoencoder (VAE) training objective. In particular, the VAE is implemented using a simple convolutional neural network architecture, the details of which are shown in Table 1. The dimensionality of latent states $|z|$ was chosen to be $16$, thus ensuring high informational bottleneck to encourage feature disentanglement. The feature extractor is trained using Adam optimiser [36] with a learning rate of $0.0001$. The batch size is $128$. A gradient step is applied after each environment step, and the training continues throughout RL training. We set the beta parameter $\beta=0.1$, which is a common way to encourage better image reconstructions in VAEs [37], and use MSE as reconstruction loss.

### v2

@@ 4.3 How sample-efficient is GoalLadder?
Using large pretrained models is expensive, so methods that rely on them for feedback must do so efficiently. While embedding-based approaches, like VLM-RM [rocamonde_vision-language_2023], use all environment observations for reward calculation, preference-based methods, like RL-VLM-F [wang_rl-vlm-f_2024], can offer better sample efficiency by training a reward function from a limited amount of feedback, in the hope that the learned reward function generalises to unseen states. By contrast, GoalLadder produces reward signals that generalise to unseen states by defining the reward in terms of distances in an embedding space (Section 3.2.3) trained with unsupervised learning. Utilising the full extent of agent’s experiences, rather than just the labelled observations, for reward definition allows GoalLadder to approximate state utility without collecting large amounts of feedback. Averaged across all Metaworld tasks, GoalLadder learns to solve them after only $\sim$4500 queries. For comparison, PEBBLE [lee_pebble_2021], a preference-based method that uses ground-truth reward preferences, achieves the same performance after attaining an average of $\sim$15000 preference labels on the same tasks. Indeed, we observe that GoalLadder boasts superior sample efficiency compared to the tested baselines. Performance curves in Figure 4 illustrate that RL-VLM-F struggles to learn the tasks using the feedback rate of GoalLadder, while VLM-RM, that embeds all collected observations, fails almost entirely.
@@ A.1 Soft-Actor Critic
The implementation of the Soft-Actor Critic (SAC) [haarnoja_soft_2018] used in the paper follows the implementation of SAC in PEBBLE [lee_pebble_2021], similar to wang_rl-vlm-f_2024 to ensure a fair comparison. As such, the used hyperparameters, including the used optimisers, are the same and can be found in the original paper [lee_pebble_2021].
@@ A.2 Feature extractor
As mentioned in the paper, we train a visual feature extractor $\psi(\cdot)$ using a variational autoencoder (VAE) training objective. In particular, the VAE is implemented using a simple convolutional neural network architecture, the details of which are shown in Table 1. The dimensionality of latent states $|z|$ was chosen to be $16$, thus ensuring high informational bottleneck to encourage feature disentanglement [zakharov_long-horizon_2023]. The feature extractor is trained using Adam optimiser [kingma_adam_2017] with a learning rate of $0.0001$. The batch size is $128$. A gradient step is applied after each environment step, and the training continues throughout RL training. We set the beta parameter $\beta=0.1$, which is a common way to encourage better image reconstructions in VAEs [higgins_beta-vae_2017, zakharov_episodic_2020, zakharov_variational_2022], and use MSE as reconstruction loss.
@@ Appendix B Task descriptions and prompts
Appendix C Additional ablations
C.1 ELO rating
To disentangle the contributions of the ELO ranking more thoroughly, we have run a small ablation study of GoalLadder’s performance without the ELO system, i.e. the ’greedy’ rating system where we always trust the VLM’s decision to replace the top-rated goal. Table 5 shows the results on the Drawer Open task, confirming the importance of the ELO rating system in achieving good performance.
Table 5: ELO ablation.
GoalLadder Final success rate
with ELO 0.97 $\pm$ 0.11
without ELO 0.20 $\pm$ 0.35
Qualitatively, we observe the expected suboptimal behaviour in GoalLadder without ELO: significantly better goals are periodically replaced by significantly worse goals as targets (i.e. VLM makes mistakes). This destabilises the training process and results in a rapid performance drop.
C.2 Buffer size
In our initial experimentation, we have found that the buffer size should be big enough to allow for diversity, but small enough to allow for frequent pairwise comparisons between the candidate goals.
Table 6: Success rates by buffer size.
Buffer size Final success rate Average success rate Max success rate
1 (No ELO) 0.20 $\pm$ 0.35 0.14 $\pm$ 0.23 0.25 $\pm$ 0.45
5 0.83 $\pm$ 0.13 0.48 $\pm$ 0.13 1.00 $\pm$ 0.00
25 0.80 $\pm$ 0.11 0.47 $\pm$ 0.14 1.00 $\pm$ 0.00
50 1.00 $\pm$ 0.00 0.61 $\pm$ 0.11 1.00 $\pm$ 0.00
10000 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01 0.13 $\pm$ 0.27
To confirm this intuition, we have run a small ablation of GoalLadder with varying buffer sizes. The results shown in Table 6 confirmed our initial hypothesis: (a) smaller buffer sizes hinder performance as the diversity of candidate goals in the buffer suffers; (b) large buffer sizes require an exponentially larger number of pairwise comparisons to arrive at converged ratings, thus also hindering the performance of the model. Overall, we find that GoalLadder is robust to buffer sizes within the reasonable ranges.
C.3 Visual encoders
We have similarly tested the performance of GoalLadder with two off-the-shelf visual encoders – DINOv2 [oquab_dinov2_2024] and CLIP [radford_learning_2021] – shown in Table 7. These results indicate that a small VAE trained on environment observations outperforms large, pre-trained encoders. We hypothesise that the VAE allows the model to better distinguish between fine-grained state differences, which are not well captured by the pre-trained models.
Table 7: Success rate by vision encoder.
Model Max success rate
GoalLadder + VAE 1.00 $\pm$ 0.00
GoalLadder + DINOv2 0.25 $\pm$ 0.21
GoalLadder + CLIP 0.38 $\pm$ 0.17